A silver coulometer was connected in series with another cell containing M²^+ solution of an unknown metal. M. After passing electric current for 1hr through the two cell it was found that 5.43g of Ag was deposited, while 2.83g of M was deposited in the cell. What is the molar mass of M?


1
Expert's answer
2022-04-16T05:34:42-0400

Faraday law for electrolysis

m = z × I × t

z - electrochemical equivalent

I = m/(z × t) = (m × n × F) / (Mr × t) = (5.43 g × 1 × 96500 C) / (107.87 g/mol × 3600 s) = 1.35 A

m(M) = z × I × t = (Mr(M) × I × t)/(n × F)

Mr(M) = (m(M) × n × F) / (I × t) = (2.83 g × 2 × 96500) / (1.35 A × 3600 s) = 112.38 g/mol